OpenStudy (anonymous):

Which of the following equations could be used to solve for the tenth term of the following sequence? 15, 13, 11, 9, ... A(10) = 15 + 10(-2) A(10) = 15 + 9(-2) A(10) = 15 + 9(2) A(10) = 15 + 10(2)

OpenStudy (yrelhan4):

it is an AP with common difference -2 and the first term 15.. nth term of an AP a(n)=a+(n-1)d = 15 + (10-1)(-2) = 15 + 9(-2)

OpenStudy (yrelhan4):

and.. whenever you like an answer, you click 'best response' :)

OpenStudy (yrelhan4):

@lunagamez12